# Geometric quantum thermodynamics: A fibre bundle approach

### v1

@@ I Introduction
The paper is structured as follows. In Sec. II, we review the gauge theory of the thermodynamics group developed in Refs. [20, 21]. Section III presents how gauge theories can be formulated in terms of the geometry of principal fibre bundles and introduces some common examples from field theory to pave the way for the construction of our theory, which is done in Sec. IV. Finally, in Section V, we discuss how our geometric theory may help in understanding thermodynamics and provide new mathematical insights into the theory by using geometrical and topological techniques, as well as proposing future directions for investigation. Additionally, in Appendix A, we revise the mathematics of fibre bundles, principal bundles, and Lie theory, which are used extensively throughout this paper, in order to make our discussion more self-contained.
@@ II Quantum thermodynamics as a gauge theory
Classical thermodynamics is a theory concerned with the macroscopic bulk properties of large systems, where fluctuations vanish. The theory focuses on the average behaviour of the microscopic degrees of freedom (justified by the central limit theorem), while ignoring the microscopic configurations. This means that, in classical thermodynamics, we have limited access to information, which is referred to as coarse graining [1].
@@
To eliminate the redundancy created by this excess information, the theory introduces the concept of an emergent gauge symmetry. This symmetry is called emergent because the excess information contained in a quantum system is redundant only from the perspective of thermodynamics and not from information theory [20]. This is analogous to the classical case, where the coarse graining is implemented by ignoring the microscopic degrees of freedom that are not accessible to classical measurements. Building on the notion of redundant information introduced in Refs. [20, 21], density matrices, which act as carriers of information, are proposed to be interpreted as analogues of gauge potentials in gauge theories.
@@
The main idea behind the theory is to identify a symmetry group such that any density operator linked to $\rho_{t}$ by the action of the group gives the same value for the mean energy. In this way, we have an equivalence class of systems such that
@@
respectively. In these equations, the integration is performed over time, considering a thermodynamic process that occurs between times 0 and $\tau$. See Refs. [20, 21] for the physical meaning of these quantities. It is important to observe here that we have two distinct actions on the system. First, we have a physical process under which the state evolves, and we have the action of the thermodynamic group, which is just a symmetry action at each instant of time. Work and heat, as well as other physical quantities, are related to the physical process.
@@
[…]
In Eq. (8), the $\Pi_{n_{t}^{k}}$ are projectors onto the subspace associated with the degeneracy $n_{t}^{k}$, while $\mathbb{I}_{n_{t}^{k}}$ stands for the identity operator on that subspace. The index $E$ stands for the energy eigenbasis. If we choose a distinct observable, the structure of this operator will be the same, but the eigenprojectors will be those of the other observable. See Ref. [21] for further details.
@@ IV Geometric theory of quantum thermodynamics
Thus, we have thus far constructed a principal $\mathrm{U}(d)$-bundle $\xi$ and defined a connection on it. It is now time to understand how physical quantities such as $\rho_{t}$ and $H_{t}$ arise in this construction. To do this, we need to build an associated vector bundle.
@@
Given the above argument, we naturally consider the density matrices not as emergent gauge potentials, as proposed in Refs. [20, 21], but as specific matter fields of the theory. Thus, in this framework, the thermodynamic quantities of interest arise as functionals of the matter fields (observables), much like in a quantum field theory several physical properties arise as functionals of matter fields (for example, Dirac or Klein-Gordon fields) [8, 9, 25].
@@
This second geometric structure defined above formalises the idea that the emergent gauge of quantum thermodynamics arises only concerning measurements, which we assume are taken at some specific time $t$. This can be understood as the geometric meaning of the coarse-graining introduced in Ref. [20]. In a situation where the degeneracies of the Hamiltonian (and consequently the thermodynamic group) do not change continuously in time but vary in specific times $\{t_{1},t_{2},...,t_{m}\}$, we can define, instead of bundles over points, bundles over time intervals $0,t_{1}),[t_{1},t_{2}),...[t_{m-1},t_{m}),[t_{m},\infty)$ equipped with connections given by the Maurer-Cartan forms and Theorem 2. This construction may be useful in a geometrical study of quench dynamics, where one turns on additional interaction terms in the Hamiltonian at a discrete set of times.
One important thing to note is that, since the curvature form is a 2-form and can also be written as the unique pullback of a 2-form in the base space [23], our principal bundles have globally vanishing curvature due to the fact that a 1-dimensional base space such as $\mathds{R}$ can have no non-vanishing 2-forms.
@@ V Discussion
Furthermore, although the structure associated with the thermodynamic group does not, in general, form a single principal fibre bundle, it may form a collection of them if the degeneracies of the observables under consideration change discretely over time. In particular, formalising these ideas geometrically allows us to put forward a series of questions regarding how the topological properties of the thermodynamic group influence the thermodynamics of a quantum system.
@@ VI Acknowledgments
LCC acknowledges CNPq through grant 308065/2022-0, the financial support of the National Institute of Science and Technology for Applied Quantum Computing through CNPq grant 408884/2024-0, and the warm hospitality of the International Institute of Physics, where this work was developed with the support of the Simons Foundation (Grant No. 1023171, R.C.).
@@ A.3 Principal Bundles
be a fundamental vector field on $\mathds{B}\times\mathrm{G}$. Then

### v2

@@ I Introduction
The paper is structured as follows. In Sec. II, we review the gauge theory of the thermodynamics group developed in Refs. [20, 21]. Section III presents how gauge theories can be formulated in terms of the geometry of principal fibre bundles and introduces some common examples from field theory to pave the way for the construction of our theory, which is done in Sec. IV. Finally, in Section V, we discuss how our geometric theory may help in understanding thermodynamics and provide new mathematical insights into the theory by using geometrical and topological techniques, as well as proposing future directions for investigation. Additionally, in Appendix A, we revise the mathematics of fibre bundles, principal bundles, and Lie theory, which are used extensively throughout this paper to make our discussion more self-contained.
@@ II Quantum thermodynamics as a gauge theory
Classical thermodynamics is a theory concerned with the macroscopic bulk properties of large systems, where fluctuations vanish. The theory focusses on the average behaviour of the microscopic degrees of freedom (justified by the central limit theorem), while ignoring the microscopic configurations. This means that, in classical thermodynamics, we have limited access to information, which is referred to as coarse graining [1].
@@
Figure 1: The coarse graining process defined by restricting the measurements one may perform defines the emergent gauge group of quantum thermodynamics through the removal of redundancies in description due to inaccessible degrees of freedom.
To eliminate the redundancy created by this excess information, the theory introduces the concept of an emergent gauge symmetry. See Fig. 1. This symmetry is called emergent because the excess information contained in a quantum system is redundant only from the perspective of thermodynamics and not from information theory [20]. This is analogous to the classical case, where the coarse graining is implemented by ignoring the microscopic degrees of freedom that are not accessible to classical measurements. Building on the notion of redundant information introduced in Refs. [20, 21], density matrices, which act as carriers of information, are proposed to be interpreted as analogues of gauge potentials in gauge theories.
@@
The main idea behind the theory is to identify a symmetry group such that any density operator linked to $\rho_{t}$ by the action of the group gives the same value for the mean energy. See Fig. 2. In this way, we have an equivalence class of systems such that
@@
Figure 2: The action of the thermodynamic group $\mathrm{G}_{T}$. The orbits (orange blobs), created by the group action $g\cdot\rho$ (which is just conjugation by the matrices $V_{t}$) are subspaces of density matrices that are indistinguishable for the measurement scheme that defines the group.
@@
respectively. In these equations, the integration is performed over time, considering a thermodynamic process that occurs between times 0 and $\tau$. See Refs. [20, 21] for the physical meaning of these quantities. It is important to note that we have two distinct actions on the system. First, we have a physical process under which the state evolves, and we have the action of the thermodynamic group, which is simply a symmetry action at each instant of time. Work and heat, as well as other physical quantities, are related to the physical process.
@@
$\rho^{E}_{t}=\bigoplus_{k=1}^{p\leq d}\frac{Tr[\Pi_{n_{t}^{k}}\rho_{t}\Pi_{n_{t}^{k}}]}{n_{t}^{k}}\mathbb{I}_{n_{t}^{k}}.$ (8)
In Eq. (8), the $\Pi_{n_{t}^{k}}$ are projectors onto the subspace associated with the degeneracy $n_{t}^{k}$, while $\mathbb{I}_{n_{t}^{k}}$ stands for the identity operator on that subspace. The index $E$ stands for the energy eigenbasis. If we choose a distinct observable, the structure of this operator will remain the same, but the eigenprojectors will be those of the other observable. See Ref. [21] for further details.
@@ IV Geometric theory of quantum thermodynamics
Thus, we have constructed a principal $\mathrm{U}(d)$-bundle $\xi$ and defined a connection on it. It is now time to understand how physical quantities such as $\rho_{t}$ and $H_{t}$ arise in this construction. To do this, we need to build an associated vector bundle.
@@
Given the above argument, we naturally consider the density matrices not as emergent gauge potentials, as proposed in Refs. [20, 21], but as specific matter fields of the theory. Thus, in this framework, the thermodynamic quantities of interest arise as functionals of the matter fields (observables), much like in a quantum field theory where several physical properties arise as functionals of matter fields (for example, Dirac or Klein-Gordon fields) [8, 9, 25].
@@
This second geometric structure defined above formalises the idea that the emergent gauge of quantum thermodynamics arises only concerning measurements, which we assume are taken at some specific time $t$. This can be understood as the geometric meaning of the coarse-graining introduced in Ref. [20]. In a situation where the degeneracies of the Hamiltonian (and consequently the thermodynamic group) do not change continuously in time but vary in specific times $\{t_{1},t_{2},...,t_{m}\}$, we can define, instead of bundles over points, bundles over time intervals $0,t_{1}),[t_{1},t_{2}),...[t_{m-1},t_{m}),[t_{m},\infty)$ equipped with connections given by the Maurer-Cartan forms and Theorem 2. This construction may be useful in a geometrical study of quench dynamics, where additional interaction terms in the Hamiltonian are turned on at a discrete set of times.
One important thing to note is that, since the curvature form is a 2-form and can also be written as the unique pullback of a 2-form in the base space [23], our principal bundles have globally vanishing curvature due to the fact that a 1-dimensional base space such as $\mathds{R}$ cannot have no non-vanishing 2-forms.
@@ V Discussion
Furthermore, although the structure associated with the thermodynamic group does not, in general, form a single principal fibre bundle, it may form a collection of them if the degeneracies of the observables under consideration change discretely over time. In particular, formalizing these ideas geometrically allows us to propose a series of questions regarding how the topological properties of the thermodynamic group influence the thermodynamics of a quantum system.
@@ VI Acknowledgments
The authors are indebted to Sebastian Deffner, Rômulo César Rougemont Pereira, Rafael Chaves Souto Araujo, and Thiago Rodrigues Oliveira for critically reading and commenting on the manuscript. LCC acknowledges CNPq through grant 308065/2022-0, the financial support of the National Institute of Science and Technology for Applied Quantum Computing through CNPq grant 408884/2024-0, and the warm hospitality of the International Institute of Physics, where this work was developed with the support of the Simons Foundation (Grant No. 1023171, R.C.).